# Scalable Multi-Objective Reinforcement Learning with Fairness Guarantees using Lorenz Dominance

### v1

@@ 6 A Large Scale Many-Objective Environment
Figure 3: Two real-world instances of the MO-TNDP environment in Xi’an (China) (?) and Amsterdam (Netherlands). (A) shows the aggregate Origin-Destination Demand per cell (sum of incoming and outgoing flows); (B) shows the group of each cell, based on the house price index quintiles.
@@
Additional directional constraints can be imposed on the agent action space. The environment code enables developers to modify the city object, incorporating adjustments to grid size, OD matrix, cell group membership, and directional constraints, making it adaptable to any city. It supports both creating transport networks from scratch and expanding existing ones. MO-TNDP is available online  [^2] Githup repository: https://github.com/dimichai/mo-tndp.
7 Experiments
@@
We focus on two cities: Xi’an in China (841 cells, 20 episode steps) and Amsterdam in the Netherlands (1645 cells, 10 episode steps). Group membership for cells is determined by the average house price, which is divided into 2-10 equally sized buckets. Figure 3 illustrates two instances of the MO-TNDP environment for five objectives (a map of group membership for all objectives is provided in the supplementary material). LCN is built using the MORL-Baselines library and the code is attached as supplementary material (?).
@@
Through a Bayesian hyperparameter search of 100 runs, we tuned the batch size, learning rate, ER buffer size, number of layers, and hidden dimension across all reported models, environments, and objective dimensions (details in the supplementary material). We compare LCN with two state-of-the-art multi-policy baselines: PCN (?) and GPI-LS (?), on widely used MO evaluation metrics. To fairly compare them, we trained all algorithms for a maximum of $30,000$ steps.
7.1 Evaluation Metrics
@@
We use three common MO evaluation metrics to compare our method with the baselines. Hypervolume: an axiomatic metric that measures the volume of a set of points relative to a specific reference point and is maximized for the Pareto front. In general, it assesses the quality of a set of non-dominated solutions, its diversity, and spread (?).
$\texttt{HV}(CS,{\bm{v}}_{ref})=\texttt{Volume}\left(\bigcup_{\pi\in CS}\left[{\bm{v}}_{ref},{\bm{v}}^{\pi}\right]\right),$ (10)
@@
where $CS$ is the set of non-dominated policies, $\texttt{Volume}(\cdot)$ computes the Lebesgue measure of the input space and $[{\bm{v}}_{ref},{\bm{v}}^{\pi}]$ is the box spanned by the reference and policy value.
Expected Utility Metric (EUM): a utility-based metric that measures the expected utility of a given set of solutions, under some distribution of utility functions. EUM measures the actual expected utility of the policies, and is more interpretable compared to axiomatic approaches like the hypervolume (?). The score is defined as:
$\texttt{EUM}(CS)=\mathbb{E}_{P_{u}}\left[\max_{\pi\in CS}u({\bm{v}}^{\pi})\right],$ (11)
@@
where $CS$ is the set of non-dominated policies, $P_{u}$ is a distribution over utility functions (we use 100 equidistant weight vectors as was done in (?)) and $\max_{\pi\in CS}u(v^{\pi})$ is the value of the best policy in the $CS$, according to the utility function $u$, defined by the sampled weights.
Sen Welfare: a welfare function, inspired by Sen’s social welfare theory, that combines total efficiency (total satisfied demand) and equality into a single measure. Equality is expressed through the Gini coefficient—a quantification of the Lorenz curve that measures reward distribution among groups (with 00 indicating perfect equality and $1$ perfect inequality) (?). The final score is measured as follows:
$\texttt{SW}(\pi)=(\sum_{i}v^{\pi}_{i})(1-\texttt{GI}({\bm{v}}^{\pi})),$ (12)
@@
where $\sum\limits_{i}\bm{v^{\pi}}_{i}$ is the sum of the returns of all objectives in policy $\pi$, and $\texttt{GI}(v^{\pi})$ is its Gini Index. We use this metric for comparative purposes, reflecting a balanced scenario where both efficiency and equality are considered. Sen welfare has been utilized in economic simulations employing RL before (?). A higher Sen welfare value signifies increased efficiency and equity.
7.2 Results
In addition to the MO-TNDP environment, we also tested our methods on the DeepSeaTreasure environment, a common two-objective MORL benchmark where an agent controls a submarine and collects treasures (?). The agent faces a trade-off between fuel consumption and treasure value — moving uses fuel, but treasures at greater depths tend to be more valuable. Since the time objective is negative, only the baseline LCN is applicable in our experiments.
@@ 7.2 Results
In Table 1, we show that LCN performs on par with PCN on the DeepSeaTreasure environment. In the rest of this section, we focus on MO-TNDP. The results discussed here are based on 5 different random seeded runs. Figure 4 (a) presents a comparison between PCN, GPI-LS and LCN across all objectives in the MO-TNDP-Xi’an environment, and Figure 4 (b) compares the vanilla LCN to the reference point alternatives.
@@
Figure 4: (a) LCN outperforms PCN & GPI-LS across all objectives in the Sen Welfare measure (Xi’an). Additionally, LCN outperforms PCN in hypervolume when the number of objectives $>4$ and in EUM for objectives $>6$, showcasing its scalability over the objective space. (b) A comparison of the trained policies of the proposed LCN, LCN-Redist and LCN-Mean models.
Table 1: Results on the DeepSeaTreasure environment.
@@
7.3 LCN outperforms PCN on many objective settings
As shown in Figure 4 (a), GPI-LS exhibits significantly lower performance across all objectives compared to PCN and LCN (GPI-LS was only tested up to $6$ objectives due to the exponentially increasing runtime). This performance gap is primarily due to the combination of the high-dimensional state-action-reward space and the limited number of timesteps. The original experiments on larger domains were conducted over more than $200$k steps; to ensure a fair comparison, we have restricted training to $30$k steps.
We focus on PCN vs LCN on the rest of the discussion. PCN shows strong performance on hypervolume and EUM in environments with 2–6 objectives. This is expected, as PCN is designed to learn diverse, non-Pareto-dominated solutions, which maximize these metrics. However, for larger objectives, LCN surpasses PCN even in hypervolume and EUM — metrics it is not specifically designed for. This occurs because, given the high state-action space of the environment, the non-Pareto-dominated solutions significantly expand with objectives, making the supervised training of PCN challenging. Furthermore, we observe that when scaling above seven objective, hypervolume for PCN collapses. In contrast, LCN effectively scales across the objective space, leveraging the smaller non-Lorenz-dominated set.
@@ 7.3 LCN outperforms PCN on many objective settings
Figure 5: Learning Curves for EUM on 3 and 10 objectives (curves for all objectives are in the supplementary material).
Figure 6: Learning Curves for Sen Welfare on 3 and 10 objectives (curves for all objectives are in the supplementary material).
7.4 Reference points improve training
@@
In Figure 4 (b), we compare the reference point mechanisms (LCN-Redist and LCN-Mean) with vanilla LCN. While the axiomatic hypervolume metric shows minimal impact of reference points on the model’s performance, the utility-based metrics, particularly EUM and Sen Welfare, reveal a different story. Vanilla LCN performs well when the objective space is limited. However, as the number of objectives increases, the introduction of reference points increases stability and outperforms the raw distance from non-dominated points.
@@
Specifically, LCN-Redist demonstrates superior performance in EUM when $d>5$. Although this seems counterintuitive, the redistribution mechanism, which creates a reference point with equal objectives, also ensures that all objectives are represented, even if they were absent in the original vector. This approach promotes policies that offer desirable trade-offs across the solution spectrum. The effectiveness of LCN-Redist is further illustrated by the learning curves in Figure 5.
Conversely, LCN-Mean performs the worst in EUM as the number of objectives increases. This outcome can be attributed to the potential skewing of the mean vector by large disparities among dimensions. If a group is consistently underrepresented in the collected experiences, its mean dimension will be close to zero, negatively affecting EUM.
@@
However, LCN-Mean performs great in Sen Welfare across all objectives, offering better stability than LCN-Redist. This result is expected, as LCN is designed to maximize Sen Welfare. LCN-Mean effectively balances outliers and creates reference points that balance efficiency and equality with minimal intervention. The learning curves for Sen Welfare are presented in Figure 6.
@@
Figure 7: $\lambda$-LCN offers flexibility between emphasizing fair distribution of rewards ($\lambda=0$) and a relaxation that allows less fair alternatives ($\lambda=1$).
@@
Figure 8: Parallel coordinate plots of the approximated front for varying values of $\lambda$. As $\lambda$ increases, the solutions become less constrained by fairness, resulting in a larger coverage set.
@@
7.5 $\lambda$-LCN can be used to achieve control over the amount of fairness preference
@@
In Figure 7, we illustrate the flexibility of $\lambda$-LCN in achieving diverse solutions across different fairness preferences. When $\lambda\approx 0$, there is little flexibility, as the goal is to find the most equally distributed policies. This results in high performance for Sen Welfare but, as expected, lower performance in terms of hypervolume due to the concentrated solutions, which cover a smaller area. On the other hand, when $\lambda\approx 1$, the space of accepted solutions expands, leading to a larger hypervolume. However, by accepting less fair solutions, the overall Sen Welfare decreases, adding an extra layer of complexity to the decision-making process. On the right side of the figure, we also show two transport lines originating from the same cell, one for $\lambda=0$ and the other for $\lambda=1$. While the lines follow a similar direction, the placement of stations and the areas they traverse can differ substantially based on the fairness preference.
@@
In Figure 8, we present parallel coordinate plots of the learned coverage sets for various values of $\lambda$. Here, too we demonstrate that as $\lambda$ increases, the flexibility for distributing rewards less fairly among different groups also increases. This leads to an expanded coverage set, providing the decision maker with more policy trade-offs.
8 Conclusion

### v2

@@ 6. A Large Scale Many-Objective Environment
Two real-world instances of the MO-TNDP environment: Xi’an, China and Amsterdam, Netherlands. Panels A and C show aggregate Origin-Destination demand per cell for Xi’an (A) and Amsterdam (C). Panels B and D show group membership of each cell based on house price index quintiles for Xi’an (B) and Amsterdam (D).
Figure 3. Two real-world instances of the MO-TNDP environment in Xi’an (China) (wei_city_2020) and Amsterdam (Netherlands). (A) and (C) show the aggregate Origin-Destination Demand per cell (sum of incoming and outgoing flows) for Xi’an (A) and Amsterdam (C). (B) and (D) show the group membership of each cell, based on the house price index quintiles for Xi’an (B) and Amsterdam (D).
@@
Given the discrete and episodic nature of this particular problem, we set the discount factor $\gamma$ to 1. The transition function $\mathbf{P}$ is deterministic, and each episode starts in the same state.
@@
Additional directional constraints can be imposed on the agent’s action space. The environment code enables developers to modify the city object, incorporating adjustments to grid size, OD matrix, cell group membership, and directional constraints, making it adaptable to any city. It supports both creating transport networks from scratch and expanding existing ones. MO-TNDP is available online [^2] Githup repository: https://github.com/dimichai/mo-tndp.
7. Experiments
To evaluate our methods and compare them against baselines, we conduct experiments in two environments. One of these is the Deep Sea Treasure environment — a widely used benchmark in Multi-Objective Reinforcement Learning (MORL) (vamplew_empirical_2011). In this environment, an agent pilots a submarine through a grid-based ocean map, aiming to collect treasures located at various depths. The agent must navigate a trade-off between two conflicting objectives: maximizing treasure value and minimizing fuel consumption. Movement consumes fuel, while treasures located deeper in the sea yield higher rewards. This setup creates a trade-off between short, low-value paths and longer, more rewarding ones. Notably, because the time objective is modeled as a negative reward (i.e., a cost), only the baseline Linear Combination of Normals (LCN) method is applicable in our experiments.
@@ 7. Experiments
In MO-TNDP, which we described in Section 6, we run experiments on two cities: Xi’an in China (841 cells, 20 episode steps) and Amsterdam in the Netherlands (1645 cells, 10 episode steps). Group membership for cells is determined by the average house price, which is divided into 2-10 equally sized buckets. Figure 3 illustrates two instances of the MO-TNDP environment for five objectives (a map of group membership for all objectives is provided in the supplementary material). LCNs are built using the MORL-Baselines library and the code is attached as supplementary material (felten_toolkit_2023).
@@
Through a Bayesian hyperparameter search of 100 runs, we tuned the batch size, learning rate, ER buffer size, number of layers, and hidden dimension across all reported models, environments, and objective dimensions (details in the supplementary material). We compare LCN with two state-of-the-art multi-policy baselines: PCN (reymond_pareto_2022) and GPI-LS (alegre_gpi), on widely used MO evaluation metrics. To fairly compare them, we trained all algorithms for a maximum of $30,000$ steps.
7.1. Evaluation Metrics
We use three common Multi-Objective (MO) evaluation metrics to compare our method with the baselines: one axiomatic metric (hypervolume) and two utility-based metrics (expected utility and Sen welfare). Axiomatic metrics evaluate the quality of a solution set by assuming that the true Pareto front represents the optimal solution. They do so without requiring any knowledge of the decision-maker’s preferences over objectives, instead focusing purely on the Pareto-optimality and diversity of the solutions (hayes_practical_2022). In contrast, utility-based metrics assume that the decision-maker has preferences over the objectives. These metrics either encode a specific utility function or assume a distribution over (or class of (zintgraf2015quality)) possible utility functions to evaluate the solutions (hayes_practical_2022).
@@ 7.1. Evaluation Metrics
Hypervolume: an axiomatic metric that measures the volume of a set of points relative to a specific reference point and is maximized for the Pareto front. In general, it assesses the quality of a set of non-dominated solutions, its diversity, and spread (hayes_practical_2022). It’s defined as:
(10) $\texttt{HV}(CS,{\bm{v}}_{ref})=\texttt{Volume}\left(\bigcup_{\pi\in CS}\left[{\bm{v}}_{ref},{\bm{v}}^{\pi}\right]\right),$
@@
where $CS$ is the set of non-dominated policies, $\texttt{Volume}(\cdot)$ computes the Lebesgue measure of the input space, and ${\bm{v}}_{ref},{\bm{v}}^{\pi}]$ is the box spanned by the reference and policy value.
Expected Utility Metric (EUM): a utility-based metric that measures the expected utility of a given set of solutions, under some assumed distribution of utility functions. Since the true utility function of the decision-maker is unknown, we sample utility functions from a prior distribution and evaluate, for each sampled utility, the maximal score within the solutions set. The EUM score is then computed as the average utility across all sampled decision-makers. Specifically, in this paper, we assume that utilities can be expressed as linear combinations of the objective values, i.e., $u({\bm{v}}^{\pi})=w^{\top}{\bm{v}}^{\pi}$, where $w$ is a non-negative weight vector representing a particular preference trade-off among objectives. For each weight vector $w$: $w_{i}\geq 0$ and $\sum_{i}w_{i}=1$, a standard weighted-sum multi-objective utility function. We sample a set of 50 well-spaced weight vectors via the Riesz s-Energy method, which ensures diversity by spreading the samples as evenly as possible over the space of valid utility weights (see (ref_dirs_energy) for details). EUM measures the actual expected utility of the policies, and is more interpretable compared to axiomatic approaches like the hypervolume (hayes_practical_2022). The score is defined as:
(11) $\texttt{EUM}(CS)=\mathbb{E}_{P_{u}}\left[\max_{\pi\in CS}u({\bm{v}}^{\pi})\right],$
@@
where $CS$ is the set of non-dominated policies, $P_{u}$ is a distribution over utility functions (we use 100 equidistant weight vectors as was done in (alegre_gpi)) and $\max_{\pi\in CS}u(v^{\pi})$ is the value of the best policy in the $CS$, according to the utility function $u$, defined by the sampled weights.
Sen Welfare: this metric is based on a welfare function, inspired by Amartya Sen’s social welfare theory. It integrates total efficiency and equality into a unified metric. Equality is quantified using the Gini coefficient, a widely used statistical measure of inequality originally developed to assess income or wealth distribution within a population. It is derived from the Lorenz curve, which plots the cumulative proportion of total reward (or resource) received by the bottom fraction of groups, ordered from the least to the most rewarded.
The Gini coefficient is defined as the area between the Lorenz curve and the line of perfect equality (a 45-degree line), normalized by the total area under the line of perfect equality. It ranges from 0 to 1, where 0 indicates perfect equality (all groups receive equal rewards), and 1 indicates maximal inequality (all rewards are concentrated in a single group) (sen_poverty_1976).
Formally, the Sen Welfare score for a policy $\pi$ is computed as:
(12) $\texttt{SW}(\pi)=\bigg(\sum_{i}{\bm{v}}^{\pi}_{i}\bigg)(1-\texttt{GI}({\bm{v}}^{\pi})),$
@@
where $\sum_{i}\bm{{\bm{v}}^{\pi}}_{i}$ is the sum of the returns of all objectives in policy $\pi$, and $\texttt{GI}({\bm{v}}^{\pi})$ is the Gini coefficient of the return vector ${\bm{v}}^{\pi}$. We use this metric for comparative purposes, reflecting a balanced scenario where both efficiency and equality are considered. Sen Welfare has been utilized in economic simulations employing RL before (zheng_ai_2022). A higher Sen Welfare value signifies increased efficiency and equity.
7.2. Results
@@
We use the Deep Sea Treasure environment as a sanity check to verify that our methods can effectively optimize in a simple, small-scale setting. In Table 1, we show that LCN performs on par with PCN on the Deep Sea Treasure environment. We next focus on the large-scale MO-TNDP environment. The results discussed here are based on five independently seeded runs. Figure 4 (a) presents a comparison between PCN, GPI-LS and LCN across all objectives in the MO-TNDP-Xi’an environment, and Figure 4 (b) compares the vanilla LCN to the reference point alternatives.
@@
Two-panel figure: (a) Performance comparison of LCN, PCN, and GPI-LS across all objectives in the Sen Welfare measure for Xi’an, showing that LCN outperforms PCN and GPI-LS, and also scales better in hypervolume and EUM as the number of objectives increases. (b) Comparison of trained policies for the proposed LCN, LCN-Redist, and LCN-Mean models.
Figure 4. (a) LCN outperforms PCN and GPI-LS across all objectives in the Sen Welfare measure (Xi’an). Additionally, LCN outperforms PCN in hypervolume when the number of objectives >4>4 and in EUM for objectives >6>6, showcasing its scalability over the objective space. (b) A comparison of the trained policies of the proposed LCN, LCN-Redist and LCN-Mean models.
Table 1. Results on the Deep Sea Treasure environment.
@@
7.3. LCN Outperforms PCN on Many-Objective Settings
As shown in Figure 4 (a), GPI-LS exhibits significantly lower performance across all objectives compared to PCN and LCN. (We limited GPI-LS to up to six objectives due to rapidly increasing runtime.) This performance gap is primarily due to the combination of the high-dimensional state-action-reward space and the limited number of time steps. The original experiments on larger domains were conducted for over 200200k steps; to ensure a fair comparison, we restricted training to 3030k steps.
We focus the remainder of the discussion on comparing PCN and LCN. PCN demonstrates strong performance on hypervolume and EUM metrics in environments with 2 to 6 objectives. This is expected, since PCN is designed to learn diverse, non-Pareto-dominated solutions that directly maximize these metrics. However, as the number of objectives increases, LCN begins to outperform PCN, even on hypervolume and EUM. This suggests that LCN is capable of learning a diverse set of policies that cover a broad region of the solution space, despite primarily targeting fair solutions.
Although this result may appear counterintuitive at first, it can be explained by the fact that the set of non-Pareto-dominated solutions grows rapidly with the number of objectives, making supervised learning significantly more difficult for PCN. Notably, when the number of objectives exceeds seven, PCN’s hypervolume performance collapses. In contrast, LCN continues to scale effectively, benefiting from the relatively smaller size of the non-Lorenz-dominated set.
@@ 7.3. LCN Outperforms PCN on Many-Objective Settings
Learning curves showing EUM performance for 3 and 10 objectives.
Figure 5. Learning Curves for EUM on 3 and 10 objectives (curves for all objectives are in the supplementary material).
Learning curves showing Sen Welfare performance for 3 and 10 objectives.
Figure 6. Learning Curves for Sen Welfare on 3 and 10 objectives (curves for all objectives are in the supplementary material).
7.4. Reference Points Improve Training
@@
In Figure 4 (b), we compare the reference point mechanisms (LCN-Redist and LCN-Mean) with vanilla LCN. While the axiomatic hypervolume metric shows minimal impact of reference points on the model’s performance, the utility-based metrics, particularly EUM and Sen Welfare, reveal a different story. Vanilla LCN performs well when the objective space is limited. However, as the number of objectives increases, introducing reference points improves stability and outperforms using raw distances from non-dominated points.
@@
Specifically, LCN-Redist demonstrates superior performance in EUM when $d>5$. Although this seems counterintuitive, the redistribution mechanism creates a reference point with equal objectives. This ensures that all objectives are represented, even if they were absent in the original vector. This approach promotes policies that achieve balanced trade-offs across the solution space. The effectiveness of LCN-Redist is further illustrated by the learning curves in Figure 5.
@@
Conversely, LCN-Mean performs the worst in EUM as the number of objectives increases. This outcome can be attributed to the potential skewing of the mean vector by large disparities among dimensions. If a group is consistently underrepresented in the collected experiences, its mean dimension will be close to zero, negatively affecting EUM. However, LCN-Mean performs great in Sen Welfare across all objectives, offering better stability than LCN-Redist. This result is expected, as LCN is designed to maximize Sen Welfare. LCN-Mean effectively balances outliers and creates reference points that balance efficiency and equality with minimal intervention. The learning curves for Sen Welfare are presented in Figure 6.
In Figure 7, we examine how the cardinality of the final, non-dominated policy sets evolves as the number of objectives increases (in the Xi’an Environment). PCN exhibits steep growth: it offers fewer than ten policies for up to three objectives, but this number rises above sixty for ten objectives. This shows, empirically, the complexity discussed in Section 5.4, i.e., that as the number of objectives grows, the PCN policy set rapidly becomes large. In contrast, LCN and LCN-Mean maintain relatively stable cardinalities across all objective counts, with the trained policy set size remaining below ten throughout. This reflects the regularizing effect of reference points in constraining the solution space. LCN-Redist displays an interesting behavior: its cardinality remains low for up to six objectives; above six objectives it LCN-Redist’s cardinality increases, approaching the size of PCN’s set of non-dominated policies. While the redistribution mechanism curbs the growth of the ER buffer in fewer objective settings, as dimensionality increases, the filtering retains more diverse experiences, which also explains why EUM remains high for LCN-Redist.
Overall, our experiments suggest the following guidance for decision-makers: if they care about obtaining many diverse policies that cover a wide range of fair trade-offs, they should choose LCN-Redist. However, if they prioritize strict fairness and are comfortable with a smaller set of policies, LCN is preferable for problems with fewer objectives, while LCN-Mean offers stable policies for higher-dimensional objective problems.
A line chart showing how the size of non-dominated policy sets changes with the number of objectives. The PCN line rises steeply as objectives increase, while LCN and LCN-Mean remain relatively stable. The LCN-Redist line stays low until six objectives, then starts increasing, approaching the trend of PCN.
Figure 7. Cardinality of the non-dominated policy sets after training, as the number of objectives increases (Xi’an environment). PCN grows rapidly, while LCN and LCN-Mean remain manageable throughout. LCN-Redist stays low up to six objectives before increasing similarly to PCN.
A plot showing the effect of varying lambda on reward distribution. Lambda ranges from 0, emphasizing fair distribution, to 1, allowing less fair alternatives.
@@
Figure 8. $\lambda$-LCN offers flexibility between emphasizing fair distribution of rewards ($\lambda=0$) and a relaxation that allows less fair alternatives ($\lambda=1$).
@@
Parallel coordinate plots of the approximated Pareto front for different values of lambda. As lambda increases, solutions become less constrained by fairness, resulting in a larger coverage set.
Figure 9. Parallel coordinate plots of the approximated front for varying values of $\lambda$. As $\lambda$ increases, the solutions become less constrained by fairness, resulting in a larger coverage set.
@@
7.5. $\lambda$-LCN Can be Used to Achieve Control over the Degree of Fairness Preference
@@
In Figure 8, we illustrate the flexibility of $\lambda$-LCN in achieving diverse solutions across different fairness preference degrees. When $\lambda\approx 0$, there is little flexibility, as the goal is to find the most equally distributed policies. This results in high performance for Sen Welfare but, as expected, lower performance in terms of hypervolume due to the concentrated solutions, which cover a smaller area. On the other hand, when $\lambda\approx 1$, the space of accepted solutions expands, leading to a larger hypervolume. However, by accepting less fair solutions, the overall Sen Welfare decreases, adding an extra layer of complexity to the decision-making process. On the right side of the figure, we also show two transport lines originating from the same cell, one for $\lambda=0$ and the other for $\lambda=1$. While the lines follow a similar direction, the placement of stations and the areas they traverse can differ substantially based on the degree of fairness preference.
@@
In Figure 9, we present parallel coordinate plots of the learned coverage sets for various values of $\lambda$. Here, too, we demonstrate that as $\lambda$ increases, the flexibility for distributing rewards less fairly among different groups also increases. This leads to an expanded coverage set, providing the decision maker with more policy trade-offs.
8. Conclusion